# CADRE: Customizable Assurance of Data Readiness in Privacy-Preserving Federated Learning

### v1

@@ IV Integration Into Existing PPFL Frameworks
In this study, we utilize the APPFL framework to demonstrate the practical application of CADRE. APPFL is an open-source framework designed to enhance privacy and security in federated learning systems. It allows researchers to implement, test, and deploy federated learning experiments across distributed clients while ensuring data privacy.
@@
APPFL consists of six key components: an aggregator, scheduler, trainer, privacy module, communicator, and compressor. These components work together to tackle challenges such as computational disparities and security concerns in distributed machine learning, while also enabling enhanced privacy protection, supporting flexible model training on decentralized data, simulating various federated learning algorithms, implementing lossy compression for efficient data transfer, and providing a highly extensible framework for customizing aggregation algorithms, server scheduling strategies, and client local trainers. The framework supports various popular synchronous and asynchronous federated learning algorithms such as FedAvg [1], FedAvgM [28], FedBuff [29], and FedCompass [30], and incorporates differential privacy techniques [31].
@@
CADRE will be integrated into the APPFL framework as an extensible module. Users can use the extensible nature of CADRE to define the metrics, rules, and remedies for a specific FL task. This allows clients to use its functions locally. This integration enables clients to evaluate data using custom metrics and apply custom remedies if the rules are not satisfied. After evaluating the data, the client agent will compile a DR report of the evaluations. These evaluations are then aggregated by the communicator within APPFL to combine the results from all clients for review. This integration demonstrates CADRE’s ease of use and versatility within existing PPFL frameworks. Figure 4 provides a visual representation of its implementation within the APPFL framework.
Figure 4: The figure illustrates the integration of CADRE within the APPFL framework.
@@
Configuring CADRE for specific FL tasks is a straightforward process that allows users to tailor its extensible functionality to meet the unique requirements of each task. The process begins with the utilization of the base CADRE module. The base CADRE module serves as a foundational template with extensible functions. By using this template, users can create a specialized CADRE module that incorporates the necessary evaluation metrics, rules, and remedies specific to their task.
@@
Once the custom CADRE module is configured, it is seamlessly integrated into the APPFL framework by uploading it. The framework is designed to accommodate such modular additions, making the integration process smooth and efficient. To activate the newly created CADRE module, users simply update the configuration file within the APPFL framework. This involves specifying the path to the custom CADRE module file that will allow the framework to recognize and utilize it appropriately. Additionally, users can pass other relevant arguments specific to the CADRE module by defining them in the configuration file. For instance, a CADRE module may require additional inputs, such as feature indices and other identifiers, for various DR-related tasks. Figure 5 illustrates an example of this configuration, showcasing the YAML-based setup used to define a custom CADRE module.
@@
Figure 5: YAML configuration for customizing a CADRE module in FL tasks, allowing users to define evaluation metrics, rules, and remedies specific to their needs.
@@
With the integration of CADRE into the APPFL framework, users gain significant advantages that aid in making informed decisions before entering the costly training phase. As data flows through the system, CADRE automatically executes defined actions, ensuring that DR issues are addressed promptly and consistently. This automation provides users with timely interventions, allowing them to focus on strategic decisions rather than manual data remediation tasks.
@@
Additionally, the comprehensive DR reports offer transparency and accountability. These reports provide users with a clear overview of the DR actions taken and allow effective assessment of DR compliance. By reviewing the detailed evaluations without exposing any raw data, while maintaining privacy and security, users can ensure that only clean and compliant data is used. Overall, this streamlined approach highlights how easily CADRE can be adapted for different FL tasks and data modalities. This concept will enhance the flexibility and effectiveness of PPFL. The documentation and code for this integration are available online [32]
@@
Ultimately, this leads to improved model performance, as AI-ready data reduces the risk of errors and noise affecting the training process. The automated nature of CADRE also supports scalability by allowing the system to efficiently handle large datasets. This allow users to make better-informed decisions, optimizing resource allocation and minimizing risks before committing to the next phases in FL.
@@ V Evaluations
In this section, we will explore the datasets, experimental setups, and custom CADRE modules used to demonstrate how we can effectively address DR related challenges across various dimensions, including data quality, privacy, fairness, and more. Since most datasets in our study do not naturally exhibit these issues, we used various data pollution techniques, which we will also discuss here. Finally, we will illustrate how our custom DR standards are achieved by utilizing the tailored metrics, rules, and remedies within the custom CADRE modules.
@@
1 Noise Management Mean magnitude of the data (image intensities or feature values) Applied remedy when the data distribution mean exceeded a threshold (e.g., $>0.37$ for MNIST). Data points with noisy indices were removed.
2 Class Imbalance Handling Class imbalance degree [33] Applied when imbalance degree $>0$. SMOTE [34] was used to oversample the minority class.
3 Duplicate Management Proportion of duplicates Applied when duplicates proportion $>0$. Duplicates were identified and removed.
@@
5 Bias Handling Statistical parity difference [35] for Adult Income dataset and representative rate difference for TCGA-BRCA dataset Applied when metric value $>0$. Stratified resampling [36] to balance sensitive groups and labels in the Adult Income dataset, while SMOTE to oversample the minority group in the TCGA-BRCA dataset.
6 Outlier Management Proportion of outliers using Inter-quartile range (IQR) method [37] Applied when outliers proportion $>0$. Outliers were clipped at IQR bounds.
@@ V-A Datasets and Experimental Setup
In this study, we utilized six diverse datasets spanning both standard benchmarks and real-world medical research. The benchmark datasets include MNIST [39], a collection of handwritten digit images widely used for image classification; CIFAR-10 [40], which comprises color images across ten classes for object recognition tasks; and Adult Income [27], a tabular dataset from the UCI repository used to predict whether an individual’s income exceeds \$50K based on census data.
@@
As discussed in section IV, we employed APPFL to integrate CADRE and conduct the experiments. We consistently used FedAvg [1] as the primary FL algorithm across all experiments. Since MNIST, CIFAR-10, and Adult Income are not inherently FL datasets, we applied non-independent and identically distributed (non-IID) partitioning to ensure data heterogeneity. For these three datasets, we partitioned the data into 10 clients per experiment and ran the experiments for 10 global epochs. On the other hand, TCGA-BRCA and IXI Tiny datasets are genuine FL datasets, already partitioned into 6 and 3 clients, respectively. As previously mentioned, the AI-READI dataset was partitioned based on the imaging device used, resulting in four clients corresponding to the four devices. CADRE operates before the actual training phase, so FL training related configurations do not impact CADRE’s execution. However, to ensure the completeness of our experiments and to validate integration in FL tasks, we reported these configurations. For AI-READI dataset, we utilized a single node with 64GB RAM, and one NVIDIA A40 GPU on the Delta supercomputer at NCSA [43]. The rest of the experiments were conducted on an Apple M2 Max MacBook Pro with 32GB unified memory.
@@ V-B Custom CADRE Modules
Module 7 uses k-anonymity level as its metric. The remedy is applied when the anonymity level is less than or equal to 1 by ensuring that each entity remains identical from at least $k-1$ others based on quasi-identifiers [38]. Quasi-identifiers are attributes that are not unique identifiers on their own but can be combined to identify individuals. For the Adult Income dataset, the quasi-identifiers were “workclass,” “race,” and “gender.” We selected these features as the quasi-identifiers because they are commonly available in public records and, when combined, could increase re-identification risk. Similarly, for the TCGA-BRCA dataset, the quasi-identifiers included demographic and self-reported characteristics such as “age_at_index,” “ethnicity_not hispanic or latino,” “ethnicity_not reported,” “race_asian,” “race_black or african american,” “race_not reported,” and “race_white.” These attributes were chosen due to their potential to link individuals across datasets and may pose privacy concerns if not anonymized.
@@ V-D Results
Since this work precedes the training phase of the FL task, it provides valuable insights for users to analyze the data and assess its readiness for training. This pre-training evaluation helps determine whether proceeding with training is worthwhile, potentially saving significant resources. While we do not report the impact on overall model performance due to not entering the training phase, existing literature suggests that many remedies applied by these modules can influence model outcomes. For instance, high-quality, noise-free data significantly enhances model performance [4, 22], while balanced class distributions can reduce bias and mitigate model drifting issues [44, 45].
However, other factors, such as achieving perfect fairness and optimal anonymity levels, may affect different aspects of model performance. A dataset with minimal statistical parity can improve model fairness [46], though it may compromise overall performance and accuracy. Similarly, as we increase the privacy budget of the data, model accuracy tend to decrease [47]. However, users might choose to prioritize data fairness, and privacy standards over model performance. Also, memory usage optimization is crucial for FL clients, as resource-constrained edge devices have limited computational and memory capacity [48]. Efficient optimization helps maintain training efficiency while preventing performance degradation. Overall, these results demonstrate that our framework can be effectively integrated into PPFL systems to meet DR-related standards before training to conserve valuable resources and funds. Moreover, the informative DR reports simplify the process for users by providing a clear understanding of the data’s condition for the FL task and setting expectations for the training phase.
@@ VI Conclusion and Future Work
In this study, we introduced a novel framework to enhance DR in PPFL systems. The framework allows users to define CADRE modules tailored to address diverse DR challenges across various downstream tasks and data modalities. By specifying custom metrics, rules, and remedies, these modules allow clients to execute processes locally and to ensure that their data meets the necessary standards while preserving privacy. CADRE generates comprehensive DR reports that include evaluations from user-defined and standard metrics, along with visualizations that provide insights into data characteristics and DR levels. It enables users to make informed decisions before training while preserving data privacy. Experiments across six diverse datasets and seven distinct DR issues demonstrate the framework’s versatility and effectiveness. The integration of the framework into the APPFL framework highlights its practical applicability. By addressing DR issues before training, our approach conserves resources and enhances model outcomes. However, there can be requirements that necessitate compromising model performance. For instance, achieving perfect fairness may affect accuracy, and increasing privacy budgets can decrease precision. These trade-offs highlight the importance of prioritizing data standards based on specific FL task requirements. Our framework allows users to set realistic expectations for training, optimize resource utilization, and lay the groundwork for reliable and equitable FL results.
For future work, we plan to expand its applicability to a broader range of datasets and explore automated methods for CADRE to further streamline the DR process. Additionally, we will investigate more computationally intensive tasks, and explore adding custom privacy-preserved modules to CADRE for user-controlled privacy protection. This will enhance the framework’s adaptability to evolving privacy standards in PPFL environments.

### v2

@@ IV Integration Into Existing PPFL Frameworks
In this study, we utilize the APPFL framework to demonstrate the practical application of CADRE. APPFL is an open-source framework designed to enhance privacy and security in FL systems. It allows researchers to implement, test, and deploy FL experiments across distributed clients while ensuring data privacy. We chose APPFL as the testbed for CADRE due to its modular and extensible architecture, which aligns well with CADRE’s design principles and integration goals. Its built-in support for differential privacy, asynchronous and synchronous training algorithms, and flexible customization of core FL components makes it a suitable and practical platform for evaluating CADRE’s capabilities in real-world PPFL scenarios.
@@
APPFL consists of six key components: an aggregator, scheduler, trainer, privacy module, communicator, and compressor. These components work together to tackle challenges such as computational disparities and security concerns in distributed machine learning, while also enabling enhanced privacy protection, supporting flexible model training on decentralized data, simulating various FL algorithms, implementing lossy compression for efficient data transfer, and providing a highly extensible framework for customizing aggregation algorithms, server scheduling strategies, and client local trainers. The framework supports various popular synchronous and asynchronous FL algorithms such as FedAvg [1], FedAvgM [28], FedBuff [29], and FedCompass [30], and incorporates differential privacy techniques [31].
@@
CADRE will be integrated into the APPFL framework as an extensible module. Administrators can use the extensible nature of CADRE to define the metrics, rules, and remedies for a specific FL task. This allows clients to use its functions locally. This integration enables clients to evaluate data using custom metrics and apply custom remedies if the rules are not satisfied. After evaluating the data, the client agent will compile a DR report of the evaluations. These evaluations are then aggregated by the communicator within APPFL to combine the results from all clients for review. This integration demonstrates CADRE’s ease of use and versatility within existing PPFL frameworks. In Figure 4, we provide an overview of its implementation within the APPFL framework.
Figure 4: An integration of CADRE in the APPFL framework.
@@
Configuring CADRE for specific FL tasks is a straightforward process that allows administrators to tailor its extensible functionality to meet the unique requirements of each task. The process begins with the utilization of the base CADRE module. The base CADRE module serves as a foundational template with extensible functions. By using this template, administrators can create a specialized CADRE module that incorporates the necessary evaluation metrics, rules, and remedies specific to their task.
@@
Once the custom CADRE module is configured, it is seamlessly integrated into the APPFL framework by uploading it. The framework is designed to accommodate such modular additions, making the integration process smooth and efficient. To activate the newly created CADRE module, administrators simply update the configuration file within the APPFL framework. This involves specifying the path to the custom CADRE module file that will allow the framework to recognize and utilize it appropriately. Additionally, administrators can pass other relevant arguments specific to the CADRE module by defining them in the configuration file. For instance, a CADRE module may require additional inputs, such as feature indices and other identifiers, for various DR-related tasks. Figure 5 illustrates an example of this configuration, showcasing the YAML-based setup used to define a custom CADRE module.
@@
Figure 5: YAML configuration for customizing a CADRE module in FL tasks, allowing administrators to define evaluation metrics, rules, and remedies specific to their needs.
@@
With the integration of CADRE into the APPFL framework, administrators gain significant advantages that aid in making informed decisions before entering the costly training phase. As data flows through the system, CADRE automatically executes defined actions, ensuring that DR issues are addressed promptly and consistently. This automation provides administrators with timely interventions, allowing them to focus on strategic decisions rather than manual data remediation tasks.
@@
Additionally, the DR reports offer transparency and accountability. These reports provide administrators with a clear overview of the DR actions taken and allow effective assessment of DR compliance. By reviewing the detailed evaluations without exposing any raw data, while maintaining privacy and security, administrators can ensure that only clean and compliant data is used. Overall, this streamlined approach highlights how easily CADRE can be adapted for different FL tasks and data modalities. This concept will enhance the flexibility and effectiveness of PPFL. The documentation and code for this integration are available as part of APPFL [32].
@@
Integration of CADRE into APPFL leads to improved model performance, as AI-ready data reduces the risk of errors and noise affecting the training process. CADRE also supports scalability by allowing the system to efficiently handle large datasets. This allows administrators to make better-informed decisions, optimizing resource allocation and minimizing risks before committing to the next phases in FL.
@@ V Evaluations
To demonstrate the effectiveness of CADRE in evaluating data quality, privacy, and fairness, we use multiple datasets, experimental setups, and custom CADRE modules. Since most of the publicly available datasets that have been used to develop FL models are relatively clean and preprocessed, we used various data pollution techniques to evaluate with CADRE. We will illustrate how our custom DR standards are achieved by utilizing the tailored metrics, rules, and remedies within the custom CADRE modules. We will present an example illustrating the performance improvement of the final FL model on the downstream task when using CADRE, compared to without using it.
@@
1 Noise Management Mean magnitude of the data (image intensities or feature values) Applied remedy when the data distribution mean exceeded a threshold (e.g., >0.37>0.37 for MNIST). Data points with noisy indices were removed.
2 Class Imbalance Handling Class imbalance degree [33] Applied when imbalance degree >0>0. SMOTE [34] was used to oversample the minority class.
3 Duplicate Management Proportion of duplicates Applied when duplicates proportion >0>0. Duplicates were identified and removed.
@@
5 Bias Handling Statistical parity difference [35] for Adult Income dataset and representative rate difference for TCGA-BRCA dataset Applied when metric value >0>0. Stratified resampling [36] to balance sensitive groups and labels in the Adult Income dataset, while SMOTE to oversample the minority group in the TCGA-BRCA dataset.
6 Outlier Management Proportion of outliers using Inter-quartile range (IQR) method [37] Applied when outliers proportion >0>0. Outliers were clipped at IQR bounds.
@@ V-A Datasets and Experimental Setup
In this study, we used six datasets spanning both standard benchmarks and those from real-world medical research. The benchmark datasets include MNIST [39], a collection of handwritten digit images widely used for image classification; CIFAR-10 [40], which comprises color images across ten classes for object recognition tasks; and Adult Income [27], a tabular dataset from the UCI repository used to predict whether an individual’s income exceeds \$50K based on census data.
@@
As discussed in section IV, we employed APPFL to integrate CADRE and conduct the experiments. We consistently used FedAvg [1] as the primary FL algorithm across all experiments. Since MNIST, CIFAR-10, and Adult Income are not inherently FL datasets, we applied non-independent and identically distributed (non-IID) partitioning to ensure data heterogeneity. For these three datasets, we partitioned the data into 10 clients per experiment and ran the experiments for 10 global epochs. On the other hand, TCGA-BRCA and IXI Tiny datasets are genuine FL datasets, already partitioned into 6 and 3 clients, respectively. As previously mentioned, the AI-READI dataset was partitioned based on the imaging device used, resulting in four clients corresponding to the four devices. CADRE operates before the actual training phase, so FL training related configurations do not impact CADRE’s execution. However, to ensure the completeness of our experiments and to validate integration in FL tasks, we reported these configurations. For the AI-READI dataset, we utilized a single node with 64GB RAM and one NVIDIA A40 GPU on the Delta supercomputer at NCSA [43]. The rest of the experiments were conducted on an Apple M2 Max MacBook Pro with 32GB unified memory.
@@ V-B Custom CADRE Modules
Module 7 uses k-anonymity level as a metric. A remedy is applied when the anonymity level is less than or equal to 1 by ensuring that each entity remains identical from at least k1k-1 others based on quasi-identifiers [38]. Quasi-identifiers are attributes that are not unique identifiers on their own but can be combined to identify individuals. For the Adult Income dataset, quasi-identifiers were “workclass,” “race,” and “gender.” We selected these as the quasi-identifiers because they are commonly available in public records and, when combined, could increase re-identification risk. Similarly, for the TCGA-BRCA dataset, the quasi-identifiers included demographic and self-reported characteristics such as “age_at_index,” “ethnicity_not hispanic or latino,” “ethnicity_not reported,” “race_asian,” “race_black or african american,” “race_not reported,” and “race_white.” These attributes were chosen due to their potential to link individuals across datasets and may pose privacy concerns if identified.
@@ V-D Results
Although this work focuses on the pre-training phase of the FL pipeline, it offers important insights into the quality and readiness of data before initiating costly training procedures. By evaluating and improving DR early on, administrators can make informed decisions about whether to proceed with training. This will ultimately help conserving computational and organizational resources.
To demonstrate the downstream impact of CADRE, we conducted an experiment using the IXI Tiny dataset from the Flamby benchmark suite. This dataset consists of 3D brain MRI scans and is commonly used for medical image segmentation tasks, where performance is typically measured using the Dice score [44]. Figure 8 presents the average Dice scores between clients during 10 rounds of FL training. The blue curve shows performance before applying the CADRE noise management module, while the green curve reflects results after CADRE was used to remove noisy indices. The shaded regions represent the standard deviation between clients to capture the variability in performance.
Figure 8: The figure presents a comparative analysis of FL performance on the Flamby IXI Tiny dataset by evaluating the impact of CADRE’s noise handling module through Dice scores over ten training rounds.
The results indicate that applying CADRE leads to consistently improved and more stable Dice scores, particularly from round 4 onward. This suggests that CADRE’s early noise-handling improves DR and reduces inter-client variability. These factors are critical for achieving better generalization in FL models. These findings align with prior research showing that high-quality, noise-free data improves model robustness and accuracy [4, 22], and that addressing other DR dimensions, such as class imbalance, can mitigate bias and reduce model drift [45, 46].
However, other factors, such as achieving perfect fairness and optimal anonymity levels, may affect different aspects of model performance. A dataset with minimal statistical parity can improve model fairness [47], though it may compromise overall performance and accuracy. Similarly, as we increase the privacy budget of the data, model accuracy tend to decrease [48]. However, administrators might choose to prioritize data fairness, and privacy standards over model performance. Also, memory usage optimization is crucial for FL clients, as resource-constrained edge devices have limited computational and memory capacity [49]. Efficient optimization helps maintain training efficiency while preventing performance degradation. Overall, these results demonstrate that our framework can be effectively integrated into PPFL systems to meet DR-related standards before training to conserve valuable resources and funds. Moreover, the informative DR reports simplify the process for administrators by providing a clear understanding of the data’s condition for the FL task and setting expectations for the training phase.
@@ VI Conclusion and Future Work
In this study, we introduced a novel framework to enhance DR in PPFL systems. The framework allows FL administrators to define CADRE modules tailored to address diverse DR challenges across various downstream tasks and data modalities. By specifying custom metrics, rules, and remedies, these modules allow clients to execute processes locally and to ensure that their data meets the necessary standards while preserving privacy. CADRE allows administrators to set realistic expectations for training, optimize resource utilization, and lay the groundwork for reliable and equitable FL results.
In our future work, we will expand CADRE’s applicability to a broader range of usecases and explore automated methods to streamline the DR process. Additionally, we will investigate computationally intensive tasks and explore adding custom privacy-preserving modules to CADRE for user-controlled privacy protection. This will enhance the adaptability of CADRE to evolving privacy standards in PPFL frameworks.